A Carnot engine has an efficiency of 0.445, and the temperature of its cold reservoir is 434 K. (a) Determine the temperature of its hot reservoir. (b) If 5610 J of heat is rejected to the cold reservoir, what amount of heat is put into the engine?
Answer #1

(a) Carnot effieciency,

η = (Thot-Tcold)/Thot

0.445=1-434/Thot

434/Thot = 0.555

Thot=781.98 K

Amount of heat is put in the engine,

Carnot effieciency,

η =(Qsupplied-Qrejected) /Qsupplied

0.445 =1-Qrejected/Qsupplied

Qrejected/Qsupplied=0.555

Qsupplied =5610/0.555 = 10108.1081 J